Triangle ABC is a right triangle. The measure of one of the acute angles is 36 degree more than the measure of the other acute angle. What is the measure of the smallest angle in triangle ABC?
Answer by edjones(7569)

Let x be the smallest acute angle.
x+(x+36)=90
2x+36=90
2x+36-36=90-36
2x=54
2x/2=54/2
x=27 degrees.
Check:
x+36=63
27+63=90
.
Ed